# B2MAPO: A Batch-by-Batch Multi-Agent Policy Optimization to Balance Performance and Efficiency

### v1

@@ 1 Introduction
Cooperative Multi-Agent Reinforcement Learning (MARL) methods have addressed numerous challenges in both virtual and real-world scenarios, such as traffic signal control[24, 17], automated freight handling[6], and autonomous driving[29, 28]. Cooperative MARL aims to train a cluster of agents to collaborate on tasks to maximize the expected cumulative rewards. The existing MARL policy optimization methods are mainly divided into two categories: simultaneous policy optimization and sequential policy optimization. Simultaneous policy optimization (Figure 1a) like MAPPO[41] typically assumes that agents’ policies are independent. All agents update their policies simultaneously and cannot predict the dynamic policy changes of other agents. Non-stationarity problems are caused by dynamic changes in the environment from the perspective of one agent due to dynamic policy improvements of other agents. Consequently, the non-stationarity problems directly result in sub-optimal performance in complex task scenarios. To address this, the Centralized Training with Decentralized Execution (CTDE) [25] paradigm has been proposed. During centralized training, agents can access to other agents’ information and the global state, while during decentralized execution, agents execute decisions independently based on their individual policies. However, the non-stationarity problem still persists, especially when agents only have partial observations, making policy optimization more challenging.
@@
Conversely, sequential policy optimization follows a sequential update scheme. The sequential update scheme allows the agent to perceive the dynamic changes produced by preceding agents’ policies, providing another perspective for analyzing the interactive dependencies of agents[9]. Sequential policy optimization includes multiple rollout and single rollout methods. RSIPA[2], the representative of multiple rollout methods(Figure 1d), effectively solves non-stationary problems by converting non-stationary multi-agent reinforcement learning problems into stationary single-agent reinforcement learning (SARL) problems[18]. Therefore, RSIPA demonstrates robust performance. However, RPISA is sample-inefficient and has low efficiency during the training and execution phase. The low sample efficiency of RPISA is because each sample rollout is only used for updating the policy of a single agent. Consequently, updating the joint policy of $n$ agents requires conducting $n$ sample rollouts, leading to the sample inefficiency problem. Then, the single-rollout methods(Figure 1c) are proposed to improve sample and inference efficiency. Heterogeneous- agent Proximal Policy Optimization (HAPPO) [19] utilizes the single rollout samples to estimate local advantages. HAPPO solves sample inefficiency by updating the policies of all agents through a single-sample rollout. While HAPPO demonstrates monotonic improvement in the joint policy, it lacks a theoretical guarantee for the policy improvement of each individual agent. This results in poor algorithm performance. Furthermore, HAPPO still has low efficiency in the training phase due to its sequential update scheme which follows a predetermined order. Recently, Agent-by-agent policy optimization (A2PO)[40] can theoretically guarantee monotonic improvement of both joint and individual policies. A2PO also uses the single rollout scheme for policy optimization. Therefore there is no sample inefficiency problem. However, due to the sequential update scheme, A2PO also exhibits lower efficiency during training. Moreover, the training time increases with the number of agents, leading to higher time waste for joint policy training in large-scale agent clusters. Due to errors in estimating the advantage function, the algorithm may still fall into a sub-optimum, and the optimality cannot be guaranteed in practice. A2PO alleviates this problem by preferentially updating agents with large advantage functions. However, relying solely on advantages to determine the update sequence may be overly simplistic, as it ignores the dynamic dependencies among agents. Furthermore, forcing sequence granularity to be at the individual agent level is an extreme design.
@@
To ensure optimal performance, the surrogate objective based on offline policy correction is proposed, which theoretically guarantees the incrementally tightened monotonic bound with the monotonic improvement of both joint and individual policies. The theoretical properties lead to Batch by Batch Multi-Agent Policy Optimization(B2MAPO), a novel sequential batch updating scheme(Figure 1d). The training efficiency of B2MAPO is higher than that of the A2PO because it updates policies in batch order rather than agent order.
@@ 2 Problem Definition
$e^{b_{k}}=b_{1}\cup b_{2}\cup\cdots\cup b_{k-1}$ is the set of preceding batches updated before batch $b_{k}$. At time step $t$, each agent $i\in b_{k}$ obtains $o_{t}^{i}\in\Omega$ through the observation function $\Omega(s_{t},i)$ and actions $\boldsymbol{a}_{t}^{j\in e^{b_{k}}}$made by preceding batch agents. Then, by taking an action $a_{t}^{i}\in A^{i}$ based on its policy $\pi^{i}(\cdot|s_{t},\boldsymbol{a}_{t}^{j\in e^{b_{k}}})$, where $s_{t}$ is approximated by the trajectory $\tau_{t}^{i}=\{o_{1}^{i},a_{1}^{i},\cdots,o_{t}^{i},a_{t}^{i}\}\in\mathbb{T}\equiv(\Omega\times\boldsymbol{A})^{*}$, forming a joint action $\boldsymbol{a}_{t}=\{a_{t}^{1},\cdots,a_{t}^{n}\}$ and the joint policy $\boldsymbol{\pi}=(\pi^{1},\cdots,\pi^{n})$. Then, the next state $s_{t+1}$ is obtained through the transition function $P(s_{t+1}|s_{t},\boldsymbol{a}_{t})$. The shared reward $r_{t}$ is obtained through the reward function $R(s_{t},\boldsymbol{a}_{t})$. The joint policy $\boldsymbol{\pi}$ induces a normalized discounted state visitation distribution $d^{\pi}$, where $d^{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}Pr(s_{t}=s|\pi)$ and $Pr(\cdot|\pi):\mathcal{S}\mapsto[0,1]$ is the state probability function under $\boldsymbol{\pi}$. We define the value function $V^{\pi}(s)=\mathbb{E}_{\tau\sim(P,\boldsymbol{\pi})}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\boldsymbol{a}_{t})|s_{0}=s]$ and advantage function $A^{\pi}(s,\boldsymbol{a})=R(s,\boldsymbol{a})+\gamma\mathbb{E}_{s^{\prime}\sim{P(\cdot|s,\boldsymbol{a})}}[V^{\pi}(s^{\prime})]-V^{\pi}(s)$. Our training objective is to learn a joint policy following batch sequence to maximize the expected cumulative return, denoted as $\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}J(\boldsymbol{\pi})=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{\tau\sim(P,\boldsymbol{\pi})}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$.
@@ 3.1 Monotonic Improvement of B2MAPO scheme
In order to improve the efficiency of the model training, we propose a Batch by Batch Multi-Agent Policy Optimization (B2MAPO) scheme, which is a universal sequential batched policy updating approach. Inspired by A2PO[40], two properties are proven to guarantee the optimal performance of B2MAPO scheme. (Proof details can be found in Appendix)
Given a set of agents $I$, in B2MAPO scheme, let $(B,\prec)$ denote the total order set of batches (sets of agents), called batch sequence, where $B\!=\!\{b_{k}|b_{k}\!\subseteq\!I\!\wedge\bigcup b_{k}\!=\!I\wedge\forall i\!\neq\!j,b_{i}\cap b_{j}\!=\!\varnothing\}$; $b_{k}$ represents the $k$-th batch which update its policy; $\prec$ is the binary relation on $B$, i.e. $b_{i}\!\prec\!b_{j}$ iff $b_{i}$ is updated before $b_{j}$. Let $\pi^{b_{k}}$ and $\hat{\pi}^{b_{k}}$ denote the joint policy and the updated joint policy of $b_{k}$, respectively. Then, $\hat{\boldsymbol{{\pi}}}^{b_{k}}=\hat{\pi}^{b_{1}}\times\cdots\times\hat{\pi}^{b_{k}}\times{\pi}^{b_{k+1}}\times\cdots\times{\pi}^{b_{|B|}}$ indicates the joint policy of $B$ after $b_{k}$ is updated. In each round of policy updating, suppose ${\boldsymbol{\pi}}$ is the joint policy before $B$ is updated, while $\hat{\boldsymbol{\pi}}$ is the joint policy after $B$ is updated. Then, the batched policy updating process can be represented by:
@@
$\mathcal{L}_{\hat{\boldsymbol{\pi}}^{b_{k\!-\!1}}}(\hat{\boldsymbol{\pi}}^{b_{k}})\!=\!\mathcal{J}(\hat{\boldsymbol{\pi}}^{b_{k\!-\!1}})\!+\!\frac{1}{1\!-\!\gamma}\mathbb{E}_{(s,\boldsymbol{a})\!\sim\!(P,\hat{\boldsymbol{\pi}}^{b_{k}})\!}[A^{\boldsymbol{\pi},\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})]$ (2)
@@
$\mathcal{G}_{\boldsymbol{\pi}}(\hat{\boldsymbol{\pi}})=\mathcal{J}(\boldsymbol{\pi})+\frac{1}{1-\gamma}\sum_{k=1}^{|B|}\mathbb{E}_{(s,\boldsymbol{a})\sim(P,\hat{\boldsymbol{\pi}}^{b_{k}})}\bigl{[}A^{\boldsymbol{\pi},\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})\bigr{]}.$ (3)
@@
Given $(B,\prec)$, let $\alpha^{b_{k}}=D_{TV}^{\max}(\pi^{b_{k}}\|\hat{\pi}^{b_{k}})$, $\epsilon\!=\!\max_{b_{k}}\epsilon^{b_{k}}$, $\epsilon^{b_{k}}=\max_{s,\boldsymbol{a}}|A^{\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})|$, and $\xi^{b_{k}}=\max_{s,\boldsymbol{a}}|A^{\boldsymbol{\pi},\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})-A^{\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})|$, then: the monotonic bound of B2MAPO is incrementally tightened, i.e.$|\mathcal{J}(\hat{\boldsymbol{\pi}})-\mathcal{G}_{\boldsymbol{\pi}}(\hat{\boldsymbol{\pi}})|$
$\!\leq\!4\epsilon\sum_{k\!=\!1}^{|B|}\alpha^{b_{k}}\!\left(\!\frac{1}{1\!-\!\gamma}\!-\!\frac{1}{1\!-\!\gamma\left(1\!-\!\sum_{j\in(B_{k}\!\cup\{\!b_{k}\!\})\!}\alpha^{j}\right)}\!\right)\!+\!\frac{\sum_{k\!=\!1}^{|B|}\xi^{{b_{k}}}}{1-\gamma}$ (4)
@@ 3.2 B2MAPO Framework
The B2MAPO framework adopts the CTDE principle which supports decision model training based on B2MAPO scheme. The structure of B2MAPO framework is illustrated in Figure 2. It is a two-layer hierarchical structure, consisting of a batch sequence generation layer(upper layer) and a batch by batch multi-agent optimization layer(lower layer). The upper layer contains two modules, batch partitioning and batch ordering, which partition agents to batches and generate batch updating order. The lower-layer includes two joint policy networks, a joint policy $\boldsymbol{\pi}$ trained by B2MAPO scheme and a derived joint policy $\boldsymbol{\pi}_{ind}$ according to CTDE-based MARL algorithms.
@@
During centralized training, at the time step $t$, the upper-layer periodically outputs batch sequence based on $s_{t}$ for future $T$ time steps. Batch by batch multi-agent policy optimization layer utilizes the batch sequence to train $\boldsymbol{\pi}$ and $\boldsymbol{\pi}_{ind}$ in parallel. $\boldsymbol{\pi}$ and $\boldsymbol{\pi}_{ind}$ receive $s_{t}$ and joint observation $\boldsymbol{o}_{t}$ to generate joint actions $\boldsymbol{a}_{t}$. $\boldsymbol{\pi}$ performs MARL training based on the reward $r_{t}$. During the execution phase, only $\boldsymbol{\pi}_{ind}$ is employed for decision-making. At each time step $t$, $\boldsymbol{\pi}_{ind}$ outputs joint actions $\boldsymbol{a}_{t}$ based on joint observations $\boldsymbol{o}_{t}$.

### v2

@@ 1 Introduction
Cooperative Multi-Agent Reinforcement Learning (MARL) methods have addressed numerous challenges in both virtual and real-world scenarios, such as traffic signal control[24, 17], automated freight handling[6], and autonomous driving[29, 28]. Cooperative MARL aims to train a cluster of agents to collaborate on tasks to maximize the expected cumulative rewards. The existing MARL policy optimization methods are mainly divided into two categories: simultaneous policy optimization and sequential policy optimization. Simultaneous policy optimization (Figure 1a) like MAPPO[41] typically assumes that agents’ policies are independent. All agents update their policies simultaneously and cannot predict the dynamic policy changes of other agents. Therefore, due to dynamic policy improvements of other agents, from the perspective of one agent, the environment changes dynamically, which causes non-stationarity problems[14]. Consequently, the non-stationarity problems directly result in sub-optimal performance in complex task scenarios. To address this, the Centralized Training with Decentralized Execution (CTDE) [25] paradigm has been proposed. During centralized training, agents can access to other agents’ information and the global state, while during decentralized execution, agents execute decisions independently based on their individual policies. However, the non-stationarity problem still persists, especially when agents only have partial observations, making policy optimization more challenging.
@@
Conversely, sequential policy optimization follows a sequential update scheme. The sequential update scheme allows the agent to perceive the dynamic changes produced by preceding agents’ policies, providing another perspective for analyzing the interactive dependencies of agents[9]. Sequential policy optimization includes multiple rollout and single rollout methods. RSIPA[2], the representative of multiple rollout methods(Figure 1d), effectively solves non-stationary problems by converting non-stationary multi-agent reinforcement learning problems into stationary single-agent reinforcement learning (SARL) problems[18]. Although RSIPA demonstrates robust performance, its sample efficiency and inference efficiency are extremely lower than that of MARL methods with simultaneous policy updating scheme. Then, the single-rollout methods(Figure 1c) are proposed to improve sample and inference efficiency. Heterogeneous-agent Proximal Policy Optimization (HAPPO) [19] utilizes the single rollout samples to estimate local advantages for policy optimization, which improves sample efficiency. Although HAPPO introduces the first multi-agent trust region method that enjoys a theoretically justified monotonic improvement guarantee, it results in poor performance due to the lack of a theoretical guarantee for individual policy improvement. Furthermore, HAPPO still has low efficiency in the training phase because of its sequential predetermined policy updating scheme. Recently, Agent-by-agent policy optimization (A2PO)[40] theoretically guarantees monotonic improvement of both joint and individual policies. However, due to the sequential update scheme, A2PO also exhibits lower efficiency during training. Moreover, the training time increases with the number of agents, leading to higher training and execution time for joint policy optimization in large-scale agent clusters. Due to errors in estimating the advantage function, A2PO may still fall into a sub-optimum. Meanwhile, the sequential updates of HAPPO and A2PO are established upon the assumption that each agent’s update does not depend on and impact the policies of others. However, this assumption, becomes inapplicable with parameter sharing, as each agent’s policy update can induce changes in the policy network parameters of other agents, thereby the optimality cannot be guaranteed in practice. A2PO alleviates this problem by a heuristic approach, which preferentially updates agents with large advantage functions. However, relying solely on advantages to determine the update sequence may be overly simplistic, since it ignores the dynamic dependencies among agents.
@@
To ensure optimal performance, the surrogate objective based on offline policy correction is proposed, which theoretically guarantees the incrementally tightened monotonic bound with the monotonic improvement of both joint and individual policies. The theoretical properties lead to Batch by Batch Multi-Agent Policy Optimization(B2MAPO), a novel sequential batch updating scheme(Figure 1d). Training efficiency of B2MAPO is higher than that of A2PO because it updates policies in batch order rather than agent order.
@@ 2 Problem Definition
At time step $t$, each agent $i$ obtains $o_{t}^{i}\!\in\!\Omega$ through the observation function $\Omega(s_{t},i)$ and takes an action $a_{t}^{i}\!\!\in\!\!A^{i}$ based on its policy $\pi^{i}(\cdot|s_{t})$, where the global state $s_{t}$ is approximated by the trajectory $\tau_{t}^{i}=\{o_{1}^{i},a_{1}^{i},\cdots,o_{t}^{i},a_{t}^{i}\}\!\in\!\mathbb{T}$. This forms the joint action $\boldsymbol{a}_{t}=\{a_{t}^{1},\cdots,a_{t}^{n}\}$ and the joint policy $\boldsymbol{\pi}=(\pi^{1},\cdots,\pi^{n})$. Then, the next state $s_{t+1}$ is obtained through the transition function $P(s_{t+1}|s_{t},\boldsymbol{a}_{t})$. The shared reward $r_{t}$ is obtained through the reward function $R(s_{t},\boldsymbol{a}_{t})$. The joint policy $\boldsymbol{\pi}$ induces a normalized discounted state visitation distribution $d^{\boldsymbol{\pi}}$, where $d^{\boldsymbol{\pi}}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}Pr(s_{t}=s|\boldsymbol{\pi})$ and $Pr(\cdot|\boldsymbol{\pi}):S\mapsto[0,1]$ is the state probability function under $\boldsymbol{\pi}$. The value function and advantage function are defined as $V^{\boldsymbol{\pi}}(s)\!=\!\mathbb{E}_{(s_{t},\boldsymbol{a}_{t})\sim(P,\boldsymbol{\pi})}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},\boldsymbol{a}_{t})|s_{0}=s]$ and $A^{\boldsymbol{\pi}}(s,\boldsymbol{a})=R(s,\boldsymbol{a})+\gamma\mathbb{E}_{s^{\prime}\sim{P(\cdot|s,\boldsymbol{a})}}[V^{\boldsymbol{\pi}}(s^{\prime})-V^{\boldsymbol{\pi}}(s)]$, respectively. The optimization objective of joint policy is to maximize the expected cumulative return, denoted as $\boldsymbol{\pi}^{*}=\arg\max_{\boldsymbol{\pi}}\mathcal{J}(\boldsymbol{\pi})=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{s_{0}\sim d^{\boldsymbol{\pi}},(s_{t},\boldsymbol{a}_{t})\sim(P,\boldsymbol{\pi}})[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$.
@@ 3.1 Monotonic Improvement of B2MAPO scheme
In order to improve the efficiency of the model training, we propose a Batch by Batch Multi-Agent Policy Optimization (B2MAPO) scheme, which is a universal sequential batched policy updating approach. Inspired by A2PO[40], two properties are proven to guarantee the optimal performance of B2MAPO scheme. (Proof details can be found in [zhang2024b2mapo])
Given a set of agents $I$, in B2MAPO scheme, let $(B,\prec)$ denote the total order set of batches (sets of agents), called batch sequence, where $B\!=\!\{b_{k}|b_{k}\!\subseteq\!I\!\wedge\bigcup b_{k}\!=\!I\wedge\forall i\!\neq\!j,b_{i}\cap b_{j}\!=\!\varnothing\}$; $b_{k}$ represents the $k$-th batch which update its policies simultaneously;
$\prec$ is the binary relation on $B$, i.e. $b_{i}\!\prec\!b_{j}$ iff $b_{i}$ is updated before $b_{j}$. Let $\pi^{b_{k}}$ and $\hat{\pi}^{b_{k}}$ denote the joint policy and the updated joint policy of $b_{k}$, respectively. Then, $\hat{\boldsymbol{{\pi}}}^{b_{k}}=\hat{\pi}^{b_{1}}\times\cdots\times\hat{\pi}^{b_{k}}\times{\pi}^{b_{k+1}}\times\cdots\times{\pi}^{b_{|B|}}$ indicates the joint policy of $B$ after $b_{k}$ is updated. In each round of policy updating, suppose ${\boldsymbol{\pi}}$ is the joint policy before $B$ is updated, while $\hat{\boldsymbol{\pi}}$ is the joint policy after $B$ is updated. Then, the batched policy updating process can be represented by:
@@
$\!\mathcal{L}_{\hat{\boldsymbol{\pi}}^{b_{k\!-\!1}}}(\hat{\boldsymbol{\pi}}^{b_{k}})\!=\!\mathcal{J}(\hat{\boldsymbol{\pi}}^{b_{k\!-\!1}})\!+\!\frac{1}{1\!-\!\gamma}\mathbb{E}_{(s,\boldsymbol{a}\!)\!\sim\!(P,\hat{\boldsymbol{\pi}}^{b_{k}}\!)\!}A^{\boldsymbol{\pi}\!,\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})\!$ (2)
@@
$\mathcal{G}_{\boldsymbol{\pi}}(\hat{\boldsymbol{\pi}})=\mathcal{J}(\boldsymbol{\pi})+\frac{1}{1-\gamma}\sum_{k=1}^{|B|}\mathbb{E}_{(s,\boldsymbol{a})\sim(P,\hat{\boldsymbol{\pi}}^{b_{k}})}A^{\boldsymbol{\pi},\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})$ (3)
@@
Given $(B,\prec)$, let $\alpha^{b_{k}}=D_{TV}^{\max}(\pi^{b_{k}}\|\hat{\pi}^{b_{k}})$, $\epsilon\!=\!\max_{b_{k}}\epsilon^{b_{k}}$, $\epsilon^{b_{k}}=\max_{s,\boldsymbol{a}}|A^{\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})|$, and $\xi^{b_{k}}=\max_{s,\boldsymbol{a}}|A^{\boldsymbol{\pi},\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})-A^{\hat{\boldsymbol{\pi}}^{b_{k-1}}}(s,\boldsymbol{a})|$, then: the monotonic bound of B2MAPO is incrementally tightened, i.e.
$|\mathcal{J}(\hat{\boldsymbol{\pi}})-\mathcal{G}_{\boldsymbol{\pi}}(\hat{\boldsymbol{\pi}})|\leq\frac{\sum_{k\!=\!1}^{|B|}\xi^{{b_{k}}}}{1-\gamma}+\enspace\enspace\enspace\enspace\enspace\enspace\enspace\enspace$
$4\epsilon\sum_{k\!=\!1}^{|B|}\alpha^{b_{k}}\!\left(\!\frac{1}{1\!-\!\gamma}\!-\!\frac{1}{1\!-\!\gamma\left(1\!-\!\sum_{j\in(B_{k}\!\cup\{\!b_{k}\!\})\!}\alpha^{j}\right)}\!\right)$ (4)
@@ 3.2 B2MAPO Framework
Based on B2MAPO scheme, a universal modulized plug-and-play B2MAPO hierarchical framework, which adopts the CTDE principle, supports convenient integration of any MARL models to fully exploit and merge their merits, including optimal performance and inference efficiency. The structure of B2MAPO framework is illustrated in Figure 2. It is a two-layer hierarchical structure, consisting of a batch sequence generation layer(upper layer) and a batch by batch policy optimization layer(lower layer). The upper layer contains two modules, batch partitioning and batch ordering, which partition agents to batches and generate batch updating order. The lower-layer includes two joint policy networks, a joint policy $\boldsymbol{\pi}$ trained by B2MAPO scheme and a derived joint policy $\boldsymbol{\pi}_{ind}$ according to CTDE-based MARL algorithms.
@@
During centralized training, at the time step $t$, the upper-layer periodically outputs batch sequence based on $s_{t}$ for future $T$ time steps. Batch by batch policy optimization layer utilizes the batch sequence to train $\boldsymbol{\pi}$ and $\boldsymbol{\pi}_{ind}$ in parallel. $\boldsymbol{\pi}$ and $\boldsymbol{\pi}_{ind}$ receive $s_{t}$ and joint observation $\boldsymbol{o}_{t}$ to generate joint actions $\boldsymbol{a}_{t}$. $\boldsymbol{\pi}$ performs MARL training based on the reward $r_{t}$. During the execution phase, only $\boldsymbol{\pi}_{ind}$ is employed for decision-making. At each time step $t$, $\boldsymbol{\pi}_{ind}$ outputs joint actions $\boldsymbol{a}_{t}$ based on joint observations $\boldsymbol{o}_{t}$.